Assume that 20% of a very common insect species in your study area is parasitized. Assume that insects are parasitized independently of each other. If you collect 10 specimens of this species, what is the probability that at least one specimen in your sample is parasitized?
Answer #1

p = 0.2

n = 10

P(at least one is parasitized) = 1 - P(no one is parasitized) = 1 - (1 - 0.2)10 = 0.8926 (ans)